# On the Fragility of Active Learners for Text Classification

### v5

@@ 1 Introduction
Within a supervised learning setup, Active Learning (AL) techniques use a Query Strategy (QS) to identify an unlabeled set of instances which is optimal in the following sense: if labelled and added to the training data, they lead to the greatest improvement in model accuracy, relative to any other same-sized set. In cases where labelling is expensive, the value proposition of AL is that it is cost-efficient compared to random sampling, and a model reaches greater accuracy with a smaller number of labelled instances.
In practice, an AL technique is selected based on the strength of prior reported results, i.e., there are no “prerequisite checks”: tests that one might perform on an unlabeled dataset, that help to select a technique suited to a problem [^1] We refer to this as the practitioner’s decision model and formalize it in §4.4.. This trust extends to related decisions such as batch and seed sizes, as well as the hyperparameters (if any) of the AL technique since there is no way to empirically pick them: to compare with random sampling, or among techniques, labels are required. But if one had labels, they wouldn’t need AL! In this sense, the AL setup is unforgiving as one needs to make the optimal choice in one shot.
This leads us to question the validity of the implicit but consequential assumption of task transfer. A related question is if it makes sense to use AL in an “Always ON” mode in a data labeling workflow; this is akin to asking if AL might perform worse than random sampling. We need to quantify both the frequency and magnitude of gains from AL, to be able to evaluate the cost of such pipeline; even simple AL techniques require a model to be evaluated over the unlabeled data pool, which can be expensive depending on the model complexity, size of the data pool and the latency allowed per AL iteration.
@@
In all this, our motivation is to not disapprove of AL as an area for research, but to motivate the inclusion of multiple practical challenges in future studies.
Contributions: Our primary contribution is a rigorous empirical analysis of the learning behavior of AL techniques over multiple text classification pipelines, that is targeted towards answering the questions asked above. Additionally, we open source an AL evaluation framework [^3] Our code is available here: https://anonymous.4open.science/r/On_the_Fragility_of_Active_Learners/README.md., to enable researchers to not only reproduce our analysis, but also to rigorously evaluate their own contributions.
@@ 2 Previous Work
Critique of AL is not new. Attenberg and Provost (2011) criticize AL for its unpredictable (for a task) warm-up times, i.e., a minimum number of labeled instances before which gains over random sampling are evident. Margatina and Aletras (2023) point out problems with AL simulations. Lüth et al. (2023) identify key issues leading to a lack of realistic AL evaluations and propose solutions that they apply to image classification. Lowell et al. (2019) study AL empirically but focus on the interesting notion of successor models, i.e., future models that would use the labeled data collected via AL using a specific model. Zhan et al. (2021) examine the empirical effectiveness of AL, but they don’t evaluate on NLP tasks. Siddhant and Lipton (2018) is an empirical study of AL effectiveness similar in spirit to ours, but they focus on deep Bayesian methods.
@@ 4.1 Configuration Space of Experiments
Figure 1: The space of experiments is shown. See §4.1 for description. All representations are produced by pre-trained models, which are ubiquitous in practice today. The lines between the boxes “Representation” and “Classifier” denote combinations that constitute our prediction pipelines. Note that RoBERTa is an end-to-end predictor, where there are no separate representation and classification steps.
@@
Figure 1 enumerates the configuration space. For further details (including hyperparameters) see §B and §E. Note that all representations used are based on pre-trained models which have grown quite popular in the past few years. For classification, we picked one each of a linear, non-linear and Deep Learning based classifier. Since batch or seed sizes are inconsistent in AL literature, e.g., DAL, REAL and CAL respectively use batch sizes of $50$, $150$, $2280$ we vary these settings as well.
@@ 4.2 Metrics and Other Settings
As shown in Figure 1 we have $\boldsymbol{350}$ unique configurations. We also execute each configuration three times in the interest of robust reporting. This gives us a a total of $350\times 3=\boldsymbol{1050}$ trials. For each AL iteration of each of these trials, we follow the due process of model selection and calibration [^6] RoBERTa is the only exception since it is naturally well-calibrated Desai and Durrett (2020)..
@@ 4.3 Notation and Terminology
Let $f$ be a function that computes the model metric of interest, e.g., F1-macro. This accepts, as parameters, the random variables $h,q,d,b,s,n$, which are defined as follows:
@@ 5.3 Effect of Prediction Pipeline vs QS
Instead we use Kendal’s W to directly measure the effect size Tomczak and Tomczak (2014). These effect sizes for the QS and pipeline parameters respectively are $0.34$ and $0.25$; the effect size here measures agreement, i.e., using different QSes produce similar results (higher agreement), relative to using different pipelines. We also built an Explainable Boosting Machine (EBM) Lou et al. (2013, 2012) on our observations, which is a form of Generalized Additive Model that takes into account pairwise interactions. The global feature importances for QS+pipeline, QS and pipeline respectively are 0.35, 0.39, 0.61 - which (a) justifies looking at the marginal effects since the importance score for QS and pipeline independently are greater than QS+pipeline, and (b) corroborates that changing pipelines has a greater impact.
@@ 6 Summary and Conclusion
The use of RoBERTa as the prediction pipeline is the only (isolated) case where we see consistent positive relative improvements. Our hypothesis as to why is that an end-to-end classifier has a more coherent view of the overall distribution, and therefore informativeness of a sample. But, obviously, we can’t discount the role that RoBERTa’s specific pre-training might play here, and further experimentation is required to disentangle their respective influences. Even in this case, we point out that (1) it provides further evidence for the argument that the QS is not the deciding factor, and (2) while positive, the gains aren’t considerable, with $\delta\approx 1$ (see Figure 3 (a)-(f)).
@@ 7 Limitations
Being an empirical work, our conclusions are tied to the algorithms and settings analyzed.
@@ 8 Acknowledgements
We thank Sashank Gummuluri for early results that indicated that the broader area of Active Learning needs a closer investigation under practical settings. Joshua Selinger reviewed multiple drafts of the paper, and proposed alternative methods of measurement, for which we are grateful to him. We also owe a debt of gratitude to Mandar Mutalikdesai for his encouragement and continued support of this project.

### v6

@@ 1 Introduction
Within a supervised learning setup, Active Learning (AL) techniques Settles (2009) use a Query Strategy (QS) to identify an unlabeled set of instances which is optimal in the following sense: if labelled and added to the training data, they lead to the greatest improvement in model accuracy, relative to any other same-sized set. In cases where labelling is expensive, the value proposition of AL is that it is cost-efficient compared to random sampling, and a model reaches greater accuracy with a smaller number of labelled instances.
In practice, an AL technique is selected based on the strength of prior reported results, i.e., there are no “prerequisite checks”: tests that one might perform on an unlabeled dataset, that help to select a technique suited to a problem [^1] We refer to this as the practitioner’s decision model and formalize it in §4.4.. This trust extends to related decisions such as batch and seed sizes, as well as the hyperparameters (if any) of the AL technique since there is no way to empirically pick them: to compare with random sampling, or among techniques, labels are required. But if one had labels, they wouldn’t need AL Attenberg and Provost (2011)! In this sense, the AL setup is unforgiving as one needs to make the optimal choice in one shot Margatina and Aletras (2023).
This leads us to ask multiple questions about the broader area. How valid is the implicit but consequential assumption of transferability? A related question is whether the focus on QSes alone is warranted - how much do the other components of a prediction pipeline affect outcomes? And finally, does it make sense to use AL at least in an “Always ON” mode in a data labeling workflow; this is akin to asking if AL might perform worse than random sampling. We need to quantify both the frequency and magnitude of gains from AL, to be able to evaluate the cost of such pipelines. This is because even simple AL techniques require a model to be evaluated over the unlabeled data pool, which can be expensive depending on the model complexity, size of the data pool and the latency allowed per AL iteration.
@@
In all this, our motivation is not to disapprove of AL as an area for research, but to motivate the inclusion of multiple practical challenges in future studies.
Contributions: Our primary contribution is a rigorous empirical analysis of the learning behavior of AL techniques over multiple text classification pipelines, that is targeted towards answering the questions asked above. Additionally, we open source an AL evaluation framework [^3] Our framework, ALchemist, is available here: https://github.com/ThuongTNguyen/ALchemist., to enable researchers to not only reproduce our analysis, but also to rigorously evaluate their own contributions.
@@ 2 Previous Work
Critique of AL is not new. Attenberg and Provost (2011) criticize AL for its unpredictable (for a task) warm-up times, i.e., a minimum number of labeled instances before which gains over random sampling are evident. Margatina and Aletras (2023) point out problems with AL simulations. Lüth et al. (2023) identify key issues leading to a lack of realistic AL evaluations and propose solutions that they apply to image classification. Lowell et al. (2019) study AL empirically but focus on the interesting notion of successor models, i.e., future models that would use the labeled data collected via AL using a specific model. Zhan et al. (2021) examine the empirical effectiveness of AL, but they don’t evaluate on NLP tasks. Siddhant and Lipton (2018) is an empirical study of AL effectiveness similar in spirit to ours, but they focus on deep Bayesian methods. Prabhu et al. (2019) study sampling biases in deep AL, but their study is limited to one prediction model - FastText.zip Joulin et al. (2016) - and considers only QSes based on uncertainty sampling.
@@ 4.1 Configuration Space of Experiments
Figure 1: The space of experiments is shown. See §4.1 for description. All representations are produced by pre-trained models, which are ubiquitous in practice today. The lines between the boxes “Representation” and “Classifier” denote combinations that constitute our prediction pipelines. Note that RoBERTa is an end-to-end predictor, where there are no separate representation and classification steps. Also note that the popular Transformer architecture Vaswani et al. (2017) is represented by RoBERTa and MPNet here.
@@
Figure 1 enumerates the configuration space. For further details (including hyperparameters) see §B and §E. Note that all representations used are based on pre-trained models which have grown quite popular in the past few years. For classification, we picked one each of a linear, non-linear and Deep Learning based classifier [^5] Although end-to-end classifiers, e.g., RoBERTa, DistilBERT Sanh et al. (2020), are popular today, we include pipelines with separate representation and classification components since they are still used where: (a) a good latency-accuracy trade-off is needed, and (b) there are multiple downstream tasks that might leverage the representation, e.g., classification, similarity-based retrieval, sentiment analysis. On a different note, the growing popularity of Retrieval Augmented Generation (RAG) Lewis et al. (2020) has re-shifted focus to the area of learning good embeddings.. Since batch or seed sizes are inconsistent in AL literature, e.g., DAL, REAL and CAL respectively use batch sizes of $50$, $150$, $2280$ we vary these settings as well.
@@ 4.2 Metrics and Other Settings
As shown in Figure 1 we have $\boldsymbol{350}$ unique configurations. We also execute each configuration three times in the interest of robust reporting. This gives us a a total of $350\times 3=\boldsymbol{1050}$ trials. For each AL iteration of each of these trials, we follow the due process of model selection [^7] Margatina and Aletras (2023) point out that this is lacking in most AL studies. This is another way the current work differentiates itself. and calibration [^8] RoBERTa is the only exception since it is naturally well-calibrated Desai and Durrett (2020)..
@@ 4.3 Notation and Terminology
Let $f$ be a function that computes the model metric of interest, e.g., F1-macro. This accepts, as parameters, the random variables [^9] Of course, in this work we consider them to only assume specifically chosen values, e.g., RF, LinSVC and RoBERTa as predictors. $h,q,d,b,s,n$, which are defined as follows:
@@ 5.3 Effect of Prediction Pipeline vs QS
Instead we use Kendal’s W to directly measure the effect size Tomczak and Tomczak (2014). These effect sizes for the QS and pipeline parameters respectively are $0.34$ and $0.25$; the effect size here measures agreement, i.e., using different QSes produce similar results (higher agreement), relative to using different pipelines. We also built an Explainable Boosting Machine (EBM) Lou et al. (2013, 2012) on our observations, which is a form of Generalized Additive Model that takes into account pairwise interactions. The global feature importance [^15] The EBM was constructed using different train-test splits, hence both mean and standard deviations across these splits are reported. for QS+pipeline, QS and pipeline respectively are $0.41\pm 0.01$, $0.42\pm 0.02$, $0.63\pm 0.01$ which (a) justifies looking at the marginal effects since the importance score for QS and pipeline independently are at least as large as QS+pipeline, and (b) corroborates that changing pipelines has a greater impact.
@@ 6 Summary and Conclusion
The use of RoBERTa as the prediction pipeline is the only (isolated) case where we see consistent positive relative improvements. Our hypothesis as to why is that an end-to-end classifier has a more coherent view of the overall distribution, and therefore informativeness of a sample. But, obviously, we can’t discount the role that RoBERTa’s specific pre-training might play here, and further experimentation is required to disentangle their respective influences. Even in this case, we point out that (1) it provides further evidence for the argument that the QS alone does not decide outcomes, and (2) while positive, the gains aren’t considerable, with $\delta\approx 1\%$ (see Figure 3 (a)-(f)).
@@ 7 Limitations
Being an empirical work, our conclusions are tied to the algorithms and settings analyzed. In particular, the experiments (a) don’t include Large Language Models, or (b) is not exhaustive wrt hyperparameters such as batch and seed set sizes; we use two settings, but please note that there is no standard way to select values for these a priori. Another aspect that is not considered here is the difference between academic vs real-world datasets, which might lead to different behaviors for a QS Margatina and Aletras (2023)
We also point out that the shortcomings of individual QSes themselves are not a limitation of this study, which maybe seen as an empirical survey with the goal of thoroughly evaluating existing QSes as-is.
@@ 8 Acknowledgements
We thank Sashank Gummuluri for early results in various practical settings. Joshua Selinger reviewed multiple drafts of the paper, and proposed alternative methods of measurement, for which we are grateful to him. We also owe a debt of gratitude to Mandar Mutalikdesai for his encouragement and continued support of this project. The authors would also like to thank reviewers from the ACL Rolling Review process, especially YZn8, who helped improve various technical aspects of the paper.